# Beyond Scaling: Measuring and Predicting the Upper Bound of Knowledge Retention in Language Model Pre-Training

### v4

@@ Abstract
The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent. Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets. In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task that directly reflects a model’s internalized knowledge without the help of external tools. We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance. In response, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models. We then develop a multi-template QA evaluation framework incorporating paraphrased question variants. Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring additional training. Experimental results show that SMI outperforms co-occurrence-based baselines, achieving $R^{2}>0.75$ on models with over one billion parameters. Theoretical analysis further suggests an upper bound of around 80% QA accuracy under optimal pre-training, reflecting intrinsic memory limitations and motivating the use of retrieval or few-shot methods in later stages.
Abstract
The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent. Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets. In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task that directly reflects a model’s internalized knowledge without the help of external tools. We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance. In response, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models. We then develop a multi-template QA evaluation framework incorporating paraphrased question variants. Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring additional training. Experimental results show that SMI outperforms co-occurrence-based baselines, achieving $R^{2}>0.75$ on models with over one billion parameters. Theoretical analysis further suggests an upper bound of around 80% QA accuracy under optimal pre-training, reflecting intrinsic memory limitations and motivating the use of retrieval or few-shot methods in later stages.
[…]
Figure 1: Illustration of the three primary challenges addressed in this work, with the third being the most crucial.
@@ 1 Introduction
The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" [25]. However, the report provides little detail on the methodology behind this prediction approach.
This technique is critical for the pre-training of large language models (LLMs). On the one hand, the cost of a single pre-training run can reach hundreds of millions of dollars, making re-pre-training a highly challenging endeavor [14, 39, 8]. The prediction technique, which allows for accurate performance forecasting on certain tasks based only on the pre-training signals, helps mitigate the resource waste. On the other hand, it can guide the construction of pre-training data that are more aligned with target tasks, optimizing resource usage [12, 21].
In this paper, we aim to predict performance in closed-book QA, a vital downstream task indicative of a model’s internal knowledge [29]. This setting is particularly well aligned with the nature of pre-training data and best reflects a model’s ability to answer fact-based questions. Meanwhile, we encounter three major challenges (Figure 1):
Challenge1: Limited access to and understanding of pre-training corpora. Currently, few open-source foundation models fully disclose their pre-training datasets. Pre-training from scratch is prohibitively expensive, requiring massive data collection and compute. For example, Meta reports that pre-training the 405B-parameter Llama 3 model used 15.6 trillion tokens and 16,000 H100 GPUs, with estimated costs over 100 million dollars [8].
Challenge2: Limitations of current evaluation methods for pre-trained models. Given the characteristics of closed-book QA tasks, we can evaluate a model’s accuracy on these tasks to assess the factual knowledge it has internalized during pretraining [29, 36]. However, most evaluation methods face challenges, such as being overly sensitive to specific in-context examples and having coarse granularity in test data segmentation [24, 33]. These issues make it difficult to accurately assess whether the pre-trained model has effectively retained specific knowledge. In particular, assessing a model’s knowledge using a single query template is unreliable [40].
Challenge3: Limitations of frequency-based metrics in predicting model performance. The model’s success on the target task relies on its ability to acquire world knowledge during pre-training, a process heavily shaped by the data [1, 37]. Previous research suggests that models tend to acquire knowledge more effectively when it is frequently encountered within the pre-training data [16, 22, 2]. However, these frequency-based methods does not take into account the specificity of knowledge, and thus only demonstrates general trends. For example, despite similar frequencies of occurrence, the model may struggle to pinpoint the location of Apple’s headquarters due to the pervasive nature of the subject. In contrast, identifying NVIDIA’s headquarters might be relatively straightforward. Furthermore, frequency-based methods do not consider the impact of model size.
Our work primarily focuses on effectively addressing these three challenges to predict model knowledge from pre-training signals, thereby estimating the model’s performance on downstream tasks without any additional training. Specifically:
First, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained LLMs. With comprehensive access to the full pre-training data, we are able to systematically trace and examine the knowledge that each model was exposed to during pre-training, enabling a detailed understanding of the relationship between pre-training data and downstream performance.
Second, inspired by the evaluation methods for fine-tuning LLMs [40], we develop a robust multimodal evaluation approach. Recognizing that the same knowledge can correspond to various questioning styles, each style allows for the design of corresponding templates. For each category of question-answering tasks, we design 20 semantically equivalent but syntactically diverse evaluation templates, enabling a more accurate and reliable assessment of the model’s knowledge retention.
Finally, we propose SMI, an information-theoretic metric that linearly associates pre-training data characteristics, model size, and QA accuracy without requiring additional training. Experimental results on 24 pre-trained models indicate that the SMI metric effectively predicts a model’s knowledge retention. Across various model sizes, ranging from 14 million to 176 billion parameters, SMI significantly outperforms the co-occurrence-based baseline. For models with more than 1 billion parameters, the coefficient of determination (R²) between SMI and QA accuracy exceeds 0.75, indicating a strong linear relationship.
Moreover, the comparative analysis across models shows diminishing returns from both scaling model size and optimizing data. As model size increases, its marginal impact on QA performance decreases, suggesting that larger capacity alone cannot guarantee continued gains. Likewise, while higher SMI values are generally associated with better accuracy, pushing SMI further becomes increasingly difficult due to its logarithmic growth. Notably, even in the largest models within each family, QA accuracy consistently plateaus around 80% when SMI is maximized. This suggests a fundamental limitation in how much factual knowledge can be internalized during pre-training, highlighting the need for external augmentation methods such as retrieval or prompting in downstream stages.
In summary, our contributions are threefold:
@@
    We propose an information theory-based method and introduce the SMI metric, which accurately predicts the knowledge retention of pre-trained LLMs using only pre-training signals, without requiring any additional training.
@@
    Across LLMs with varying sizes from 14 million to 176 billion parameters, the SMI metric exhibits a strong linear relationship with QA accuracy, significantly outperforms the traditional co-occurrence-based baseline.
@@
    Experimental results across diverse models reveal that both scaling model size and optimizing data yield diminishing returns. Even under optimal conditions, QA accuracy plateaus around 80%, reflecting the intrinsic limit of knowledge that can be memorized during pre-training.
@@ 2 Related Work
The characteristics of pre-training data are crucial for knowledge retention and downstream performance in LLMs. Many studies emphasize the role of repetition frequency. For example, Carlini et al. [6] identified a logarithmic relationship between repetition and memory effects, while Chowdhery et al. [7] showed that sequences repeated over 500 times are completed with over 40% accuracy. Ju et al. [15] found that high-frequency data can lead to “factual shortcuts” in multi-hop reasoning. Allen-Zhu and Li [2] suggested that exposing knowledge 1000 times enables a storage capacity of two bits per parameter. Additionally, Razeghi et al. [28] and Yadlowsky et al. [38] linked lower-order co-occurrences to improved numerical reasoning, while McCoy et al. [23] associated task prevalence in pre-training with better performance on tasks like ROT13.
@@
Furthermore, the model size significantly influences the knowledge retention of LLMs. Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern [17].  Kandpal et al. [16] evaluate the BLOOM model families [30] on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters. Additionally, the memory capacity of a model closely correlates with its size [19].
@@
Knowledge triples (subject, relation, object) serve as an effective tool for evaluating the knowledge retention of LLMs. Petroni et al. [27] employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks. Subsequently, He et al. [13], Ju et al. [15] further explore the application of knowledge triples in reasoning tasks. Moreover, Wang et al. [37] and Allen-Zhu and Li [1] use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
@@ 3 Methodology
To address the three challenges outlined in Section 1, we propose a method to predict model knowledge solely from pre-training signals. We begin with prior work and explore the use of the co-occurrence metric. Inspired by information theory, we propose the Mutual Information (MI) metric that simultaneously considers both the frequency and specificity of knowledge, significantly enhancing prediction outcomes. Finally, we propose the SMI metric, a composite representation of pre-training data characteristics and model size.
@@ 3.1 Starting from the Co-occurrence Metric
Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns [16, 22, 2]. Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach [16]. There is a linear relationship between QA accuracy and the logarithm of the number of documents [16]. The logarithm of the document count serves as the co-occurrence metric, which constitutes our baseline.
@@
In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object [37]. For a given LLM, denoted as $F$, we define $F$ as mastering the knowledge triple $t$, thus demonstrating competence in the corresponding QA task, as follows:
@@
Evaluating the knowledge retention of LLMs presents challenges due to the diverse range of queries formed by the subject and relation of knowledge triples. Inspired by the evaluation method for fine-tuned LLMs [40], we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
@@
On the one hand, we calculate metrics from the pre-training data. Each QA task consists of a subject and an object. We perform document-level retrieval on the pre-training data and compute three values: $P(s)$ represents the proportion of documents containing the subject out of the total number of documents $\mathbf{N}$ in the corpus, $P(o)$ denotes the proportion of documents containing the object out of $\mathbf{N}$, and $P(s,o)$ signifies the proportion of documents containing both the subject and the object out of $\mathbf{N}$. The co-occurrence metric is given by $\log_{2}(P(s,o))$.
@@ 3.2 Considering the Knowledge Specificity to Obtain the MI Metric
The co-occurrence metric does not account for the specificity of knowledge. When entities of knowledge are common phrases that frequently appear in pre-training data, their specificity is low, making the knowledge more difficult to retain. Conversely, if the entities appear infrequently in pre-training data, their specificity is high, and the knowledge is easier to retain. For example, the address of NVIDIA’s headquarters is easier to remember than that of Apple’s headquarters, because the word "Apple" occurs more frequently across diverse contexts. MI in information theory considers both the co-occurrence of two variables and their individual specificities [31]. Therefore, we aim to use MI to improve the co-occurrence metric.
@@
[…]
@@
$MI(s,o)$ $=\log_{2}(I(s,o)).$ (3)
In the above formula, $P(s)$ and $P(o)$ penalize the number of documents containing the subject and object, respectively, to emphasize the specificity of $s$ and $o$. The term $I(s,o)$ represents the amount of shared information between $s$ and $o$, and it can also indicate the reduction in the amount of information about $o$ given $s$. During inference with LLMs, this amount of information refers to the probability of the model generating the object given the subject. Considering the uneven distribution of $I(s,o)$ (Figure 2), we take its logarithm to obtain the MI metric.

### v5

@@ Abstract
The GPT-4 technical report suggests that downstream performance can be predicted from pre-training signals, but offers little methodological detail on how to quantify this. This work address this gap by modeling knowledge retention, the capacity of a pre-trained language model to memorize factual information from its corpus, and introduce a principled method to estimate it prior to training. We propose Size-dependent Mutual Information (SMI), an information-theoretic predictor that integrates knowledge frequency, knowledge specificity, and model size to forecast closed-book question answering (QA) accuracy. SMI is validated through large-scale document retrieval over the disclosed pre-training corpora of 21 public and 3 custom models, combined with a robust multi-template QA evaluation. Experiments show that SMI significantly outperforms repetition-based baselines and achieves R² > 0.7 in predicting QA accuracy for models above 1B parameters, without additional training. The analysis further reveals diminishing returns from scaling data and model size and provides evidence for an intrinsic upper bound on knowledge retention achievable by pre-training alone, motivating retrieval and other augmentation strategies. The dataset and code are available at https://github.com/yuhui1038/SMI.
Beyond Scaling: Measuring and Predicting the Upper Bound of Knowledge Retention in Language Model Pre-Training
Changhao Jiangthanks: Equal contribution, Ming Zhang11footnotemark: 1, Yifei Cao11footnotemark: 1, Junjie Ye11footnotemark: 1, Xiaoran Fan11footnotemark: 1, Shihan Dou, Zhiheng Xi, Jiajun Sun, Yi Dong, Yujiong Shen, Jingqi Tong, Baoyu Fan, Tao Gui, Qi Zhangthanks: Corresponding author, Xuanjing Huang Fudan University, Shanghai, China IEIT Systems Co. Ltd, Shenzhen, China chjiang25@m.fudan.edu.cn, mingzhang23@fudan.edu.cn, qz@fudan.edu.cn
Figure 1: Overview of our method. The framework consists of three stages: (1) Corpus Analysis and Retrieval, where subject and object (question and answer) entities are extracted from pre-training data and their frequencies and co-occurrences are computed; (2) Information-Theoretic Metric Computation, where knowledge frequency, knowledge specificity, and model size are integrated into the size-dependent mutual information (SMI) metric; and (3) Evaluation via Closed-Book QA, where pre-trained language models are assessed using multi-template evaluation, and SMI is compared with empirical QA accuracy to analyze linear relationships and scaling effects.
@@ 1 Introduction
The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" (OpenAI, 2023). Understanding and predicting the generalization ability of large language models (LLMs) has become a key research question in the study of scaling laws (Kaplan et al., 2020). Although prior work has established empirical relationships between loss, data size, and model parameters, these results primarily offer coarse-grained predictions of overall performance and do not address what specific factual knowledge is retained from pre-training corpora. Recent studies have shown that LLMs can memorize and regurgitate training data (Carlini et al., 2021; Biderman et al., 2023a), but these findings are largely descriptive and do not provide a systematic way to predict retention for new model–corpus pairs. Developing principled approaches for this purpose is crucial both for improving our theoretical understanding of model training dynamics and for guiding practical decisions on data curation and compute allocation (Hoffmann et al., 2022; Alabdulmohsin et al., 2022; Ruan et al., ).
This work addresses this gap by modeling knowledge retention, the capacity of a pre-trained language model to memorize factual information from its corpus, and introduces a principled method to reliably estimate it prior to training. Rather than focusing only on aggregate loss curves or perplexity, we take a more fine-grained and systematic view: we ask, given a particular fact in the pre-training corpus, what is the probability that a model will be able to accurately reproduce it in a closed-book setting after training (Petroni et al., 2019)? We term this capacity knowledge retention, and formalize it as a quantitative property of a model–corpus pair. This framing allows us to reason about model memory at the level of individual facts and categories, and to make meaningful predictions before committing to the cost of full-scale training.
We propose Size-dependent Mutual Information (SMI), an information-theoretic predictor that integrates knowledge frequency, knowledge specificity, and model size to forecast closed-book question answering (QA) accuracy (Roberts et al., 2020). Since pre-trained language models are optimized only for next-token prediction rather than explicit knowledge retrieval, closed-book QA provides a natural and challenging probe for assessing factual recall. Our formulation builds on the intuition that both the frequency of a fact in the corpus and its rarity influence its likelihood of being learned, and that these effects are jointly influenced by model size. Concretely, SMI captures the mutual information between knowledge subjects and objects (or questions and answers), combining these statistics with model size. The resulting predictor is simple to compute from pre-training data and scales gracefully with corpus size, making it suitable for practical use in large-scale settings.
SMI is validated through large-scale document retrieval over the disclosed pre-training corpora of 21 public and 3 custom models, combined with a robust multi-template QA evaluation. Our experimental setup reconstructs the exposure of facts in the training data, then evaluates whether models can correctly answer corresponding questions in a closed-book setting. We design multi-template QA prompts to reduce prompt variance and report results across a diverse set of model families, covering a range from 14M to 176B parameters.
Experiments show that SMI significantly outperforms repetition-based baselines and achieves the coefficient of determination (R²) > 0.7 in predicting QA accuracy for models above 1B parameters, without additional training. Relative to heuristics based solely on repetition counts, SMI yields a stronger correlation with observed QA accuracy, particularly in the large-model regime where memorization behavior appears more stable. Moreover, we observe that SMI maintains comparable predictive performance across multiple model families and domains, indicating that it captures a generalizable aspect of factual knowledge acquisition in pre-trained language models.
The analysis further reveals diminishing returns from scaling both data and model size, suggesting that pre-training alone faces an intrinsic limit in its ability to retain factual knowledge. Recent studies have similarly observed a performance ceiling on knowledge-intensive tasks, even with substantial scaling of parameters and data (Zhang et al., 2025). These observations point to inherent limitations of purely parametric approaches: beyond a certain scale, further increases in model size or training data yield only marginal gains in closed-book performance. This motivates exploring hybrid approaches that augment parametric memory with non-parametric mechanisms, including retrieval-augmented generation, external memory modules, or targeted fine-tuning, to mitigate the ceiling effects observed in pre-training. Overall, our contributions are threefold:
@@
    We formalize knowledge retention as a quantitative metric defined over model–corpus pairs, providing a new way to analyze what models actually memorize during pre-training.
@@
    We propose SMI, a data-driven and training-free predictor that integrates knowledge frequency, knowledge specificity, and model size to estimate closed-book QA accuracy.
@@
    We empirically evaluate SMI on 24 models, showing that it achieves strong predictive accuracy, characterizing diminishing returns from scaling data and model size, and suggesting the presence of an upper bound on knowledge retention achievable by pre-training alone.
@@ 2 Related Work
The characteristics of pre-training data are crucial for knowledge retention and downstream performance in LLMs. Many studies emphasize the role of repetition frequency. For example, Carlini et al. (2023) identified a logarithmic relationship between repetition and memory effects, while Chowdhery et al. (2023) showed that sequences repeated over 500 times are completed with over 40% accuracy. Ju et al. (2024) found that high-frequency data can lead to “factual shortcuts” in multi-hop reasoning. Allen-Zhu and Li (2024b) suggested that exposing knowledge 1000 times enables a storage capacity of two bits per parameter. Additionally, Razeghi et al. (2022) and Yadlowsky et al. (2023) linked lower-order co-occurrences to improved numerical reasoning, while McCoy et al. (2023) associated task prevalence in pre-training with better performance on tasks like ROT13.
@@
Furthermore, the model size significantly influences the knowledge retention of LLMs. Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern (Kaplan et al., 2020).  Kandpal et al. (2023) evaluate the BLOOM model families (Scao et al., 2022) on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters. Additionally, the memory capacity of a model closely correlates with its size (Li et al., 2024).
@@
Knowledge triples (subject, relation, object) serve as an effective tool for evaluating the knowledge retention of LLMs. Petroni et al. (2019) employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks. Subsequently, He et al. (2024); Ju et al. (2024) further explore the application of knowledge triples in reasoning tasks. Moreover, Wang et al. (2024) and Allen-Zhu and Li (2024a) use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
@@ 3 Methodology
To predict a model’s ability to recall factual knowledge before training, we develop a framework that estimates closed-book QA accuracy directly from pre-training data statistics and model size. We first revisit prior work and evaluate a simple repetition-based co-occurrence metric as a baseline. Building on information-theoretic principles, we then introduce the Mutual Information (MI) metric, which accounts for both the frequency and specificity of knowledge items and yields improved predictive performance. Finally, we extend MI to SMI, which incorporates model size as a scaling factor, providing a more accurate and scalable predictor of knowledge retention.
@@ 3.1 Starting from the Co-occurrence Metric
Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns (Kandpal et al., 2023; Mallen et al., 2023; Allen-Zhu and Li, 2024b). Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach (Kandpal et al., 2023). There is a linear relationship between QA accuracy and the logarithm of the number of documents (Carlini et al., 2023; Kandpal et al., 2023). The logarithm of the document count serves as the co-occurrence metric, which constitutes our baseline.
@@
In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object (Wang et al., 2024). For a given LLM, denoted as $F$, we define $F$ as mastering the knowledge triple $t$, thus demonstrating competence in the corresponding QA task, as follows:
@@
Evaluating the knowledge retention of LLMs presents challenges due to the diverse range of queries formed by the subject and relation of knowledge triples. Inspired by the evaluation method for fine-tuned LLMs (Ye et al., 2024), we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
@@
On the one hand, we calculate metrics from the pre-training data. Each QA task consists of a subject and an object. We perform document-level retrieval on the pre-training data and compute three values: $P(s)$ represents the proportion of documents containing the subject out of the total number of documents $\mathbf{N}$ in the corpus, $P(o)$ denotes the proportion of documents containing the object out of $\mathbf{N}$, and $P(s,o)$ signifies the proportion of documents containing both the subject and the object out of $\mathbf{N}$. The co-occurrence metric is defined as $\log(P(s,o))$, representing the logarithm of the fraction of documents in which both the subject and object appear. For consistency with information-theoretic conventions (Shannon, 1948a), all logarithms in this paper use base 2. Since specifying the subject and object typically determines the relation, we omit the relation component for simplicity.
@@ 3.2 Considering the Knowledge Specificity to Obtain the MI Metric
The co-occurrence metric does not account for the specificity of knowledge. When entities of knowledge are common phrases that frequently appear in pre-training data, their specificity is low, making the knowledge more difficult to retain. Conversely, if the entities appear infrequently in pre-training data, their specificity is high, and the knowledge is easier to retain. For example, the address of NVIDIA’s headquarters is easier to remember than that of Apple’s headquarters, because the word "Apple" occurs more frequently across diverse contexts. MI in information theory considers both the co-occurrence of two variables and their individual specificities (Shannon, 1948b). Therefore, we aim to use MI to improve the co-occurrence metric.
@@
Unlike the MI in information theory, which considers all possible values of random variables, the MI metric within pre-training data considers only those instances in which the relevant entities actually appear. Given pre-training data $\mathbf{P}$ consisting of $\mathbf{N}$ paragraphs, a knowledge triple $(s,r,o)$, and corresponding $P(s)$, $P(o)$, and $P(s,o)$, the MI metric between $s$ and $o$ is defined as:
@@
$MI(s,o)$ $=Norm(\log(I(s,o))).$ (3)
In the above formula, $P(s)$ and $P(o)$ penalize the number of documents containing the subject and object, respectively, to emphasize the specificity of $s$ and $o$. The term $I(s,o)$ represents the amount of shared information between $s$ and $o$, and it can also indicate the reduction in the amount of information about $o$ given $s$. During inference with LLMs, this amount of information refers to the probability of the model generating the object given the subject. Considering the uneven distribution of $I(s,o)$ (see Figure 2), we apply a logarithmic transformation followed by normalization to map its values to the range $[0,1]$, resulting in the MI metric. In rare cases where $I(s,o)$ becomes zero or negative due to data sparsity or noise, such instances are regarded as anomalies and excluded from further analysis.